Three forces P 1 =i-j+k, P 2 =-i+2j-k, P 3 =j-k are acting on an object at A. What is the work done in displacing it from A to B ? (OA= 4i-3j-2k, OB= 6i+j-3k) Options: 8, 6, 4, 9

Three forces P1=i-j+k, P2=-i+2j-k, P3=j-k are acting on an object at A. What is the work done in displacing it from A to B ? (OA= 4i-3j-2k, OB= 6i+j-3k)
Options: 8, 6, 4, 9

Grade:12

1 Answers

Khimraj
6 years ago
Net force on object F = P1 +P2 + P3 = 2j -k
Displacement vector D = AB = OB -OA = 2i+4j-k
Work done will be equal to dot product of F and D
So W = F.D = 0+8+1 =9.
Hope it clears. If still you have doubt you can ask me.
If you like answer then please approve it.